### v1

Abstract
Monitoring maternal and fetal health during pregnancy is crucial for preventing adverse outcomes. While tests such as ultrasound scans offer high accuracy, they can be costly and inconvenient. Telehealth and more accessible body shape information provide pregnant women with a convenient way to monitor their health. This study explores the potential of 3D body scan data, captured during the 18-24 gestational weeks, to predict adverse pregnancy outcomes and estimate clinical parameters. We developed a novel algorithm with two parallel streams which are used for extract body shape features: one for supervised learning to extract sequential abdominal circumference information, and another for unsupervised learning to extract global shape descriptors, alongside a branch for demographic data. Our results indicate that 3D body shape can assist in predicting preterm labor, gestational diabetes mellitus (GDM), gestational hypertension (GH), and in estimating fetal weight. Compared to other machine learning models, our algorithm achieved the best performance, with prediction accuracies exceeding 88% and fetal weight estimation accuracy of 76.74% within a 10% error margin, outperforming conventional anthropometric methods by 22.22%.
keywords:
Pregnancy outcomes, 3D body scan, Machine learning
1 Introduction
Systematic prenatal care has been recognized as essential for reducing maternal and neonatal morbidity and mortality rates since the early twentieth century [1, 2]. However, some studies concluded that many in-person prenatal care visits are unnecessary [3, 4], and the prenatal care regimen has seen limited improvements despite decades of advances in diagnostic and communication technologies [2, 5]. Recent studies have suggested that incorporating telehealth modalities can improve prenatal care by reducing unnecessary trips to hospital and interventions [6, 7, 8, 9]. For pregnant women in rural communities, prenatal care through telehealth is even more crucial, as it helps overcome barriers to accessing equitable healthcare resources [6].
Although telehealth offers convenience, it has disadvantage of lacking in laboratory tests [10, 11]. To address this issue, it is crucial to explore biomarkers which can be easily obtained through widely-available devices for health monitoring. The 3D body shape is an ideal data modality that can be collected in non-invasive ways, especially in the field of obstetrics, where distinct body shape changes occur on pregnant women. Meanwhile, previous studies have shown a strong correlation between maternal anthropometric measurements and the health status of mothers and fetuses [12, 13, 14]. Ay et al. found that maternal body mass index (BMI) during pregnancy positively correlates with the fetal weight. Maternal height, pre-pregnancy BMI, and gestational weight gain are also found to correlate with potential risks of having a small or large for gestational age child [15]. Risk of preeclampsia and cesarean delivery have been associated with maternal weight, height and body circumferences [16, 17]. These promising results indicate the potential for using body shape features as additional resources in telehealth for pregnancy. Moreover, comparing to traditional anthropometric features, detailed 3D body scan data can provide more body shape information, such as the shape, volume and position of gravid uterus.
The development of 3D optical scanning technology has made it possible to capture 3D body shape data using commercial 3D body scanners [18]. Besides, some commodity smartphone apps allow users to accurately scan themselves at home by using built-in LiDAR or camera with 3D reconstruction algorithms [19]. With the accessibility and accuracy of 3D body models, there is an increased interest in developing algorithms capable of efficiently extracting information from detailed body shapes for obstetric analysis.
Currently, 3D body scanning technology in the field of obstetrics and gynecology is still in its early stages. However, a growing number of research has demonstrated the efficacy of assessing health conditions by analyzing 3D body shapes. Machine learning algorithms, in particular, have shown high potential in extracting latent information from 3D body shapes. For example, Lu et al. proposed a method to predict body fat percentage from 3D body scans using machine learning algorithms [20]. Zheng et al. designed a new shape descriptor and a neural network to predict appendicular skeletal muscle mass [21]. Wang et al. designed a bi-channel network to detect hepatic steatosis using 3D body shape reconstructed from CT iso-surfaces [22]. These studies present the possibility of employing 3D body shape in the field of obstetrics.
In this study, we aim to apply 3D body shape analysis in maternal and fetal health status assessment. Our main contributions are summarized as follows:
  1. 1.
    We used 3D body scans and basic demographic information to predict the risk of preterm labor, GDM, GH, and the likelihood of delivery by cesarean section. We also used the same body shape features to estimate fetal weight and MVP. To the best of our knowledge, this is the first study that uses 3D body scans of pregnant women to assess maternal and fetal health status.
  2. 2.
    We designed a novel hybrid neural network comprising a supervised learning stream and an unsupervised learning stream to extract shape features from 3D shape data, which is robust even on imbalanced small-sized data.
  3. 3.
    We conducted experiments comparing our algorithm using 3D body scans as inputs with the baseline method using support vector machine (SVM) and anthropomorphic measurements. Additionally, we also compared our algorithm with other well-performing machine learning algorithms using the same 3D body shape inputs. These experimental results not only indicate the feasibility of using 3D body scan to assess maternal and fetal health status, but also demonstrate the accuracy and efficiency of our new approach.
2 Related work
2.1 Association between anthropometric measurement and pregnancy health status
To simply predict the likelihood of adverse pregnancy outcomes, researchers explored the relationships between maternal anthropometric parameters and a range of adverse outcomes, and they found a strong correlation between them. Boucher et al. conducted a study which revealed associations between the probability of cesarean delivery and anthropometric measurements such as weight, BMI, waist circumference and skinfold thickness [17]. Sina et al. found that increasing BMI, weight, waist circumference, and waist-to-height ratio were associated with an increased risk of GDM, with BMI and waist-to-height ratio showing stronger correlations [23]. Ebrahimi-Mameghani et al. concluded that early pregnancy BMI and waist circumference were associated with the risk of GH, preeclampsia, and preterm labor [24].
Besides investigation of the relationship between maternal anthropometry and pregnancy outcomes, some studies also showed the associations between maternal body shape and intrauterine parameters. In the absence of ultrasound equipment, clinicians can estimate fetal weight by using formula such as Johnson’s formula, Insler and Bernstein’s formula for a series of clinical maternal measurements such as symphysis-fundal height (SFH) and abdominal circumference (AC) [25, 26, 14, 27, 28, 29]. Anggraini et al. proposed a model to estimate fetal weight based solely on fundal height (FH) [30]. Given single value as input, this method is relatively effective in detecting abnormal fetal growth, however, this model is still constrained by limited amount of information and the requirement for manual measurement. Recently, another study used deep neural networks to estimate fetal weights with multiple variables as inputs, including blood laboratory tests and medical history. The analysis result shows that pre-pregnancy weight and BMI have great impact on estimated fetal weight (EFW) [31].
These collective findings underscore the potential of utilizing body shape features to predict the likelihood of adverse pregnancy outcomes and estimate intrauterine parameters such as fetal weight, which can benefit applications in telehealth.
2.2 Applications of 3D body shape in obstetrics field and advanced methodologies for extracting information from 3D body shape
Anthropometric techniques have been widely used to capture a small part of body shape features for a long time, but these methods cannot fully represent the whole body shape [32, 33]. Moreover, although these methods do not require complex equipment, they still need to be carried out by professionals.
Recently, some researchers investigated the applications of 3D body scanning in obstetrics. Glinkowski et al. used 3D surface topography method to reveal postural changes of pregnant women and studied the relationship between spinal curvatures and low back pain during pregnancy [34]. Dathan-Stumpf et al. captured anthropometric parameters from 3D scan models of pregnant women to predict successful vaginal breech deliveries. The results showed that the prediction accuracy of the method using 3D scanning measurements is at least as good as Magnetic Resonance Imaging (MRI) diagnosis [35]. These studies further indicated the usability of 3D body scans in obstetrics. However, they are still limited by the extraction of a limited number of anthropometric measurements from 3D models rather than learning information directly from the 3D models as a new imaging modality.
To fully leverage information from the complicated representations of 3D models, which are often comprised of thousands or millions of polygons or 3D points [36], researchers are investigating more efficient body representations and methodologies and evaluating them on different medical applications. Xie et al. proposed to use frontal whole-body silhouettes for estimating body composition. Test results on estimating fat mass index and fat-free mass index supported the usability of this silhouette-based method [37]. Ng et al. proposed a method to estimate regional body composition by using regional circumferences, areas, and volumes obtained from 3D body models with a linear model [18]. Lu et al. used level circumferences to represent 3D body shapes. With the extracted level circumferences, they employed a Bayesian network to predict pixel-level body composition and body fat percentage for the 2D projection of 3D body [38]. Wang et al. constructed a body shape descriptor called Shape Map from body shape contour of CT scan slices and used it to predict whole-body fat percentage and visceral fat percentage [39]. Zheng et al. used abdominal level circumferences sampled from 3D body shapes to assess hepatic steatosis [40]. Roy et al. obtained the trajectory of the vertebral column from body shape and derived potential lateral deviations of the spine and rotation of the vertebrae. This method performed well in the estimations of the lateral deviation of the spine for mild and moderate scoliosis [41]. Su et al. explored the diagnosis of type II diabetes using features selected from 3D scan data through machine learning algorithms like backpropagation neural networks and decision trees [42]. Although these studies are not specifically conducted in obstetrics field, the various methods explored to represent 3D body shape and extract body shape features are worth learning.
3 Dataset and pre-process
3.1 Dataset
Our dataset comprises a total of 60 study participants, who were recruited at the George Washington University Medical Faculty Associates. The demographic characteristics of this study population is listed in Table 1. To mitigate the influence of confounding factors, we implemented the recruitment following these exclusion criteria: (1) being under the age of 18; (2) having carried multiple gestations; (3) having been diagnosed with an enlarged fibroid uterus; (4) having a BMI exceeding 60; (5) having any unstable medical or emotional condition or chronic disease that would preclude study participation; (6) having undergone body shape altering operation such as liposuction or plastic surgery. This data collection was approved by the Institutional Review Board (IRB).
Table 1: Demographic characteristics of study population.
Demographics N=60
Age(yr) $31.88\pm 6.04$
Race
   white 24
   African American 31
   others 5
Height(m) $1.62\pm 0.06$
Weight(kg) $71.91\pm 13.82$
Gestational age(d) $148.02\pm 8.84$
00footnotetext: Notes: For terms with continuous values, we provide mean±$\pm$±SD; for terms with discrete values, we provide number of participants.
Each participant underwent optical 3D body scans between the 18th and 24th weeks of pregnancy with Fit3D optical scanner (Fit3D, San Francisco, CA). The 3D scanning process was repeated two to three times to ensure the reliability of the collected data. The participants were required to wear fitting clothes to ensure accurate representation of their body shapes. To investigate the usability of smartphone scans, we also collected 3D body scans using smartphone with the application Polycam (Polycam Inc, San Francisco, CA), as shown in Figure 1 below. However, the precision of these smartphone scanned model needs to be tested by additional experiments. Since this study only discusses the theoretical usability of three-dimensional models for obstetric predictions and estimations, we have currently chosen this commercial Fit3D optical scanner with whose precision tested in other studies [43, 18]. The results will be used to determine the feasibility of models obtained through smartphone scanning for future telehealth applications.
Table 2: Clinical parameters of study population
Target
EFW(g) $378.40\pm 71.92$ (N=43)
MVP(cm) $4.90\pm 0.97$ (N=40)
Length of gestation(d) $268.69\pm 17.33$ (N=42)
Participants with cesarean section delivery 14 (N=42)
Participants with preterm labor 7 (N=42)
Participants with GDM 8 (N=51)
Participants with GH 8 (N=60)
00footnotetext: Notes: $N$ denotes the number of participants who have corresponding medical records. Since we have 6 participants transferred to other hospitals, 12 not delivered yet, 17 participants not having EFW value, 20 not having MVP value on their report, the number $N$ varies in different tasks. For terms including EFW, MVP and length of gestation, we provide mean±$\pm$±SD; for terms with binary results, we provide number of positive participants.
In addition to 3D scan data, we collected corresponding gestational age (GA) and basic demographic information including height and weight. To explore which adverse pregnancy outcomes and parameters are associated with body shape, we also collected the following information: EFW, MVP, indications of GDM and GH, length of gestation, and delivery type. The detail of these clinical parameters are listed in Table 2. Anthropometric measurements automatically generated by Fit3D scanner are used as the input for a SVM algorithm, which serves as the baseline method for our experiments. For each task, subjects who had incomplete information or those whose 3D scans were all unsuccessful were excluded from subsequent analysis.
Figure 1: Comparison of Fit3D scanned model and Polycam scanned model captured simultaneously.
3.2 Pre-process
Based on the anthropomorphic variables from our previous work [23, 24, 25, 26, 14, 27, 28, 29], we have identified weight, height, waist circumference, and hip circumference as having potentially strong correlations with maternal and fetal health status. Therefore, after we extracted heights and weights from participant’s medical records, we focused primarily on the abdominal body shape in the region from the lowest point of the pubic bone to the bottom of the breasts, as shown in Figure 2. For each 3D model, we manually labeled the lower and upper boundaries of this abdominal region. Following this, we performed a uniform sampling of 64 level circumferences within this region to create a sequence of the 64 measurements. This approach allows for a simplified yet relatively more informative representation of the 3D body shape by emphasizing the critical region pertaining to pregnancy health status.
Figure 2: Extracting abdominal level circumference sequence.
4 Method
Our algorithm is designed to leverage both the maternal body shape and basic demographic information for assessing maternal and fetal health status. Therefore, this algorithm employs a dual-branch structure that processes the two modalities of data separately, as depicted in Figure 3. The first branch is used to extract body shape features from 3D scan data, while the other branch is used to process the demographic features. In the first branch, we designed two parallel processing streams to comprehensively learn the 3D shape information from different perspectives and extract complementary information. The first stream of this branch is used to learn spatial dependency of the abdominal level circumference through a supervised learning algorithm, Recurrent Neural Network (RNN), which is sensitive to the local pattern between neighboring elements. The second stream is used to extract the most representative global features of the same region via an unsupervised Principal Component Analysis (PCA) algorithm. This algorithm has been successfully used in other studies based on 3D body shape analysis, which can guarantee the basic performance of the algorithm and enhance the robustness given a small dataset [44, 38, 40]. The features extracted from these two streams can complement each other and enhance the accuracy and stability of the overall algorithm. The final layer fuses the outputs of these two streams along with processed demographic information, predicting the probabilities of adverse outcomes or estimating parameters of interest.
In the subsequent subsections, we will give a detailed description of these structures.
Figure 3: Architecture of the proposed algorithm: Here, the 64 level circumferences and basic demographic information are used as input for the network. The two-stream body shape analysis branch, which is designed to process the level circumferences, consists of a supervised RNN and an unsupervised PCA.
4.1 Sequence-dependent feature extraction stream
After pre-processing, the complex 3D body model is represented by a one-dimensional sequence of level circumferences. This simplified representation can reduce the redundancy of the raw data and mitigate overfitting given small data size. While a back propagation (BP) neural network can extract abstract features from this body shape data, it alone may overlook latent spatial dependencies between sequence elements.
Therefore, to maximize the information derived from the level circumference sequence, we implemented an Elman RNN consisting of a single RNN layer. The network contains 64 steps, corresponding to the number of elements in the sequence. The 64 circumference values are sequentially input into the RNN unit based on their sequence index. Figure 4 illustrates the calculation at the $i$th time step
Unlike a typical neural network unit, the RNN unit processes not only the current input, but also considers the output from the previous operation. As depicted in Figure 4, the RNN unit processing the element with index $i$ uses the current level circumference value $x_{i}$ and the previous output $h_{i-1}$ as inputs. The two inputs are given different trainable weights $W_{x}$ and $W_{h}$. After adding a trainable bias $b$, the weighted sum is processed through a nonlinear activation function $f(*)$. In our network, we specifically use the hyperbolic tangent function $tanh(*)$ to implement this nonlinear transformation. The output $h_{i}$ is then sent back to the RNN cell in the next step. The entire computation process can be formulated as follows:
$h_{i}=tanh(x_{i}W^{T}_{x}+h_{i-1}W^{T}_{h}+b).$ (1)
This output is subsequently used in the next step, forming a recurrent operation. This operation facilitates the extraction of local sequential information and its conveyance towards the final output. The output of the last step $h_{out}$ will serve as the body shape feature with spatial information, encapsulating both sequence element values and spatial dependency.
Figure 4: Processing details of the $i$th element.
4.2 Global feature extraction stream
In the global feature extraction stream, we ignore the spatial relationship among level circumferences, with focus instead placed on extracting the most representative features based on their values. Unsupervised learning algorithms are usually used to discover the hidden patterns of data without label guidance. They can explore the data based on its own distribution. In our method, we employ PCA to extract the features for downstream tasks. This algorithm identifies directions in high dimensional space which explain a maximum amount of variance. In other words, it reduces the feature dimensionality while retaining most of the information.
Since smaller dataset with higher data dimensionality can increase the risk of overfitting, we use PCA to transform the redundant 64-dimensional level circumferences to lower-dimensional global features. By calculating the cumulative explained variance, we discovered that the first three principal components account for 98.1% of the variance.
Consequently, we use this 3-dimensional feature, $vec_{PCA}$, in subsequent processes as the global feature of the 3D body shape.
4.3 Joint processing layer
In addition to the 3D body shape data, demographic features also provide crucial information for the prediction and estimation tasks. In our method, we incorporate height, weight and GA as input. Our final joint processing layer concatenates the two body shape feature vectors - $h_{out}$ derived from the sequence-dependent feature extraction stream, and $vec_{PCA}$ from the global feature extraction stream - with the demographic information vector $vec_{basic}$, to make up the fused feature $vec_{fused}$.
To limit the elements from $h_{out}$ to be at the same scale as the elements from the other two normalized vectors, we used batch normalization operation before the concatenation and obtain the $vec_{RNN}$.
$vec_{fused}=concat(vec_{RNN},vec_{PCA},vec_{basic}).$ (2)
A linear fully connected layer is then applied to generate the final output. The implementation of this operation is represented by the following formula, where $W_{L}$ and $b_{L}$ are trainable network parameters:
$output=vec_{fused}W^{T}_{L}+b_{L}.$ (3)
Notice here both the dimensions of $vec_{PCA}$ and $vec_{basic}$ are 3. While processing concatenated features with fully connected layer, the dimension of each feature can also be considered as a “weight” [45]. Thus, we set the dimension of $h_{out}$ as 5, which is closer to the lower dimension of other features but still able to contain enough information, so that none of these three features can dominantly contribute to the output. In our experiments, we make minor adjustment on the final layer for different tasks. For regression tasks, we directly use this $output$ value as the final result and use the Mean Square Error (MSE) loss as criterion. For binary classification tasks, we generate a probability ranging from 0 to1 by using sigmoid function on the $output$ and use corresponding Binary Cross-Entropy (BCE) loss as criterion.
5 Experiments
We conducted binary classification tasks using 3D body shape features for the prediction of various adverse pregnancy outcomes, including the risk of preterm labor, GDM, GH, and the likelihood of undergoing a cesarean section during delivery. We also conducted regression tasks to estimate current fetal weight and MVP with the same inputs.
To assess the effectiveness of employing 3D body shape in these tasks and evaluate the performances of different algorithms, we implemented different evaluation metrics for classification tasks and regression tasks. Additionally, we designed an ablation study to validate the necessity and impact of critical components of our algorithm.
5.1 Evaluation metrics
In evaluating the binary classification performance of algorithms, we use metrics including accuracy, precision (also referred to as positive predictive value, PPV), recall, specificity, F1 score, and area under the receiver operating characteristic curve (AUC-ROC). For AUC-ROC score, the Mann-Whitney U-test is employed as a statistical test. With this method, we also calculate the p-value of the AUC-ROC score to show the effectiveness of the algorithms.
To evaluate the regression performance of algorithms, we use Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), Mean Absolute Percent Error (MAPE, also referred to as Mean Relative Error, MRE), and Root Mean Squared Percent Error (RMSPE) as the evaluation metrics.
$MAE=\frac{1}{n}\sum_{n=1}^{n}|\hat{y}_{i}-y_{i}|.$ (4)
$RMSE=\sqrt{\frac{1}{n}\sum_{n=1}^{n}(\hat{y}_{i}-y_{i})^{2}}.$ (5)
$MAPE=\frac{1}{n}\sum_{n=1}^{n}|\frac{\hat{y}_{i}-y_{i}}{y_{i}}|\times 100\%.$ (6)
$RMSPE=\sqrt{\frac{1}{n}\sum_{n=1}^{n}(\frac{\hat{y}_{i}-y_{i}}{y_{i}})^{2}\times 100\%}.$ (7)
Additionally, we introduce an accuracy metrics which gives the percentage of estimated values falling within an acceptable range. The accuracy metric is defined as follows:
$acc=\frac{1}{n}\sum_{n=1}^{n}bool(|\frac{\hat{y}_{i}-y_{i}}{y_{i}}|<m).$ (8)
Here, $m$ represents an error tolerance range. We set $m=10\%$ and $m=5\%$ respectively, which is considered acceptable for clinicians [46]. Corresponding results were calculated in our experiments.
5.2 Classification Performance
We conducted binary classification experiments to predict the risks of preterm labor, GDM, GH, and the likelihood of undergoing a cesarean section. To evaluate the effectiveness of using 3D body scans compared to traditional anthropometric measurements, we developed a baseline method using anthropometric measurements automatically generated by the optical scanner (height, weight, BMI, waist circumference, hip circumference, waist-to-hip ratio, waist-to-height ratio) and GA as inputs. Support Vector Classification (SVC) was used as the baseline model to process these inputs. We also conducted comparison experiments with several popular machine learning algorithms using the same 3D body shape and demographic features as our algorithm used. These algorithms include Logistic Regression (LoR), BP Neural Network (BPNN), Random Forest (RF), and SVC. Since these algorithms cannot effectively process high dimensional features given small data size, we incorporate PCA to help improve their performance and mitigate potential overfitting problem. A 5-fold cross-validation was used in the experiments to ensure the algorithms’ generalizability and to provide more reliable comparisons.
Table 3: Performance of classification methods for delivery type prediction
Accuracy Precision Recall Specificity F1 Score AUC-ROC (p-value)
Baseline 42.86% 36.11% 92.86% 17.86% 0.52 0.531 (0.7458)
PCA+LoR 54.76% 35.29% 42.86% 60.71% 0.39 $<0.5$ (N/A)
PCA+BPNN 57.14% 30.00% 21.43% 75.00% 0.25 $<0.5$ (N/A)
PCA+RF 54.76% 36.84% 50.00% 57.14% 0.42 0.551 (0.5937)
PCA+SVC 42.86% 14.28% 14.28% 57.14% 0.14 $<0.5$ (N/A)
Ours 28.57% 25.00% 57.14% 14.28% 0.35 $<0.5$ (N/A)
Notes: The p-value in parentheses is calculated using the Mann-Whitney U test on the AUC-ROC.
Table 4: Performance of classification methods for preterm labor prediction
Accuracy Precision Recall Specificity F1 Score AUC-ROC (p-value)
Baseline 80.95% 33.33% 14.28% 94.28% 0.20 0.592 (0.4486)
PCA+LoR 71.42% 22.22% 28.57% 80.00% 0.25 $<0.5$ (N/A)
PCA+BPNN 78.57% 37.5% 42.86% 85.71% 0.40 $<0.5$ (N/A)
PCA+RF 69.05% 25.00% 42.86% 62.86% 0.26 0.594 (0.4370)
PCA+SVC 83.33% 50.00% 28.57% 94.28% 0.36 0.539 (0.7471)
Ours 88.10% 75.00% 42.86% 97.14% 0.55 0.682 (0.1323)
Notes: The p-value in parentheses is calculated using the Mann-Whitney U test on the AUC-ROC.
Table 5: Performance of classification methods for GDM prediction
Accuracy Precision Recall Specificity F1 Score AUC-ROC (p-value)
Baseline 68.00% 21.43% 37.50% 73.81% 0.27 0.613 (0.3140)
PCA+LoR 76.47% 30.00% 37.50% 83.72% 0.33 0.703 (0.0705)
PCA+BPNN 84.31% 50.00% 37.50% 93.02% 0.43 0.743 (0.0304)
PCA+RF 86.27% 66.67% 25.00% 97.67% 0.36 0.695 (0.0823)
PCA+SVC 84.31% 50.00% 37.50% 93.02% 0.43 0.637 (0.2222)
Ours 88.24% 75.00% 37.50% 97.67% 0.50 0.843 (0.0022)
Notes: The p-value in parentheses is calculated using the Mann-Whitney U test on the AUC-ROC.
Table 6: Performance of classification methods for GH prediction
Accuracy Precision Recall Specificity F1 Score AUC-ROC (p-value)
Baseline 78.33% 14.29% 12.50% 88.46% 0.13 $<0.5$ (N/A)
PCA+LoR 83.33% 41.67% 62.50% 86.54% 0.50 0.827 (0.0031)
PCA+BPNN 85.00% 40.00% 25.00% 94.23% 0.31 0.669 (0.1263)
PCA+RF 83.33% 41.67% 62.50% 86.54% 0.50 0.867 (0.0009)
PCA+SVC 88.33% 55.56% 62.50% 92.31% 0.56 0.776 (0.0125)
Ours 88.33% 60.00% 37.50% 96.15% 0.46 0.738 (0.0313)
Notes: The p-value in parentheses is calculated using the Mann-Whitney U test on the AUC-ROC.
The results are shown in Table 6, 6, 6, 6. We considered the situations with adverse outcomes as the positive class. For instance, in the delivery type classification task, participants who underwent a cesarean section are labeled as positive.
Figure 5: Visualization of raw 64 level circumferences data. The colors and line styles are assigned according to classification results of our algorithm in different tasks.
From these results, we can conclude that 3D body shapes captured in the second trimester can effectively be used to predict risk of preterm labor, GDM and GH, which can achieve accuracy rates exceeding 88%. However, using this data cannot reliably predict the delivery type. Even with the most effective algorithm for this task, the Random Forest, the accuracy only reaches 54.76%, with a precision of 36.84%, recall rate of 50%, and AUC-ROC of 0.551. This could be due to the complex factors influencing delivery type such as previous delivery experiences and umbilical cord status, which were not included in this study.
By comparing the performances of the baseline method (SVC using anthropometric measurements as input) and the SVC using 3D body shape data, we can conclude that the same algorithm SVC can yields better prediction performance with 3D body shape as inputs than with traditional anthropometric measurements. In comparison to other algorithms, our algorithm shows superior performance in predicting risk of preterm labor, GDM and GH. Especially in predicting the risk of preterm labor and GDM, our algorithm surpasses others across all metrics. In the GH prediction task, our algorithm performs best in accuracy, precision, and specificity, while LoR, RF and SVC achieve a higher recall score of 62.5%. For predicting both GDM and GH tasks, our algorithm achieves high AUC-ROC scores with p-values lower than 0.05. These results suggest that 3D body shape data has higher potential to be used in predicting adverse pregnancy outcomes than anthropometric measurements. And these results also verified the effectiveness of our design in exploring useful information from 3D body scans.
To provide a more intuitive understanding of the relationship between abdominal body shape and adverse pregnancy outcomes, in Figure 5, the raw 64 abdominal level circumferences of subjects are presented as curve lines, and each curve is classified by colors and line styles according to the classification results yielded by our algorithm. In these charts, curves of negative subjects are drawn with blue lines and those of positive subjects are drawn with red lines. The curves of correctly classified subjects are drawn with solid lines and those of misclassified subjects are drawn with dashed lines. By observing the color of lines, we notice that the data is imbalanced, especially for preterm labor, GDM and GH. Despite the imbalance and small size of the dataset, our method yields better results in predicting preterm labor, GDM, and GH. Most negative samples are correctly classified, with false positive rates below 5%. Additionally, the details of the graph reveal intriguing information about the association between adverse outcomes and body shape patterns. We find that positive samples of cesarean section delivery do not have distinctive distribution. This may be one reason why we cannot accurately predict the cesarean section with 3D body shape and limited demographic information. Preterm labor and GDM exhibit similar positive sample distributions, which more likely happens to pregnant women with larger body circumferences. While positive GH samples tend to have relatively larger gravid uterus but smaller hip-region circumferences. From the chart we also notice some sudden changes in circumference values. This may be caused by scanner error or posture changes of participants.
5.3 Regression Performance
Besides predicting the risk of adverse outcomes, we also examined the applications of 3D body shape in estimating current intrauterine parameters. We adjusted the final layer of our network and conducted regression experiments to estimate fetal weight and MVP. In these experiments, we designed a similar baseline method using Support Vector Regression (SVR) with traditional anthropometric measurements as inputs. We also included the average of the ground truth values as an additional baseline. Linear regression (LR), BPNN, RF, and SVR with PCA were tested as comparisons methods , using the same inputs as in the classification tasks. The evaluation was conducted using 5-fold cross-validation, consistent with the approach in the classification experiments. The results are presented in Table 8 and Table 8 below.
Table 7: Performance of models for EFW estimation
MAE(g) RMSE(g) MAPE RMSPE Acc(m=10%) Acc(m=5%)
Baseline 34.31 47.57 9.33% 13.36% 62.79% 39.53%
Average 52.04 71.25 13.71% 17.93% 55.81% 20.93%
PCA+LR 41.54 56.42 11.39% 15.73% 55.81% 39.53%
PCA+BPNN 39.40 55.84 10.94% 16.05% 67.44% 34.88%
PCA+RF 43.26 60.69 11.64% 16.60% 62.79% 41.86%
PCA+SVR 38.44 48.27 10.46% 13.62% 60.46% 30.23%
Ours 30.41 43.62 8.28% 12.31% 76.74% 53.49%
Table 8: Performance of models for MVP estimation
MAE(cm) RMSE(cm) MAPE RMSPE Acc(m=10%) Acc(m=5%)
Baseline 0.786 0.988 16.18% 20.20% 37.5% 22.5%
Average 0.747 0.929 15.68% 19.85% 42.50% 27.50%
PCA+LR 1.248 1.552 25.02% 29.79% 20.00% 15.00%
PCA+BPNN 1.126 1.848 23.57% 29.32% 27.50% 12.50%
PCA+RF 0.838 1.091 17.60% 22.97% 45.00% 27.50%
PCA+SVR 0.830 1.035 17.12% 21.16% 40.00% 20.00%
Ours 0.976 1.137 20.29% 24.28% 17.50% 12.50%
Figure 6: Regression results of EFW.
The results from Table 8 indicate that our algorithm outperforms other algorithms in estimating EFW across all metrics. While SVR also demonstrates commendable performance, our algorithm exceeds SVR by 20.89% in MAE, 9.63% in RMSE, 20.84% in MAPE, 9.62% in RMSPE, and achieves an accuracy rate of 76.74% within a 10% error tolerance range. Interestingly, the baseline method also yields promising results, outperforming the same SVR algorithm with 3D body shape as inputs. Figure 6 illustrates the estimation result of each participant calculated by different algorithms. For clarity, the participants were reordered by increasing order of their actual EFW values collected from ultrasound report. From Figure 6 we find that except the average line, all other estimation results follow the trend of the actual value, with our method yielding results that fluctuate the least around the ground truth. These findings align with the observations from Table 8, establishing the utility of 3D body shape in estimating fetal weight.
When we attempted to estimate MVP value using these algorithms, the performances were not as good as when estimating fetal weights, as shown in Table 8. The MAPEs and RMSPEs of machine learning methods on this task exceed 15% and 20% respectively, which is notably higher than the roughly 10% MAPEs and 15% RMSPEs when estimating fetal weight. The RF algorithm achieves the highest accuracies, but none of these machine learning methods surpass the estimation using average value of the ground truth. This suggests that 3D body shape may not be an effective tool for estimating MVP. The line chart with reordered subjects is provided in Figure 7. In contrast to the results in Figure 6, we can barely find the relevance among the distributions of these estimation results, and none appears to align with the increasing trend of the actual value.
Figure 7: Regression results of MVP.
5.4 Ablation Study
Our proposed method exhibits overall good performance in both classification and regression tasks, which may be attributed to the separate application of PCA and RNN to extract complementary features. To substantiate this claim, we conducted a set of experiments that estimate fetal weight using networks with or without either of these two streams. Additionally, we explored replacing the RNN with the gated recurrent unit (GRU), an enhanced version of the vanilla RNN, to extract sequential information from level circumferences. The GRU is designed with ability of learning both long-term and short-term dependencies, while having fewer parameters than the long short-term memory (LSTM) network. Through comparisons with this GRU network, we aim to demonstrate that our network architecture performs better with the conventional RNN. Table 9 lists the results of the various combinations we tested during the algorithm development process.
Table 9: Performances on estimating fetal weight with different network architectures. The ”FC” denotes the fully connected layer.
MAE RMSE MAPE RMSPE Acc(m=10%)
FC 43.68 56.46 11.80% 15.47% 48.84%
PCA+FC 39.40 55.84 10.94% 16.05% 67.44%
RNN+FC 35.31 49.50 9.74% 14.26% 60.46%
PCA+GRU+FC 31.95 45.77 8.79% 13.16% 72.09%
Ours
(PCA+RNN+FC) 30.41 43.62 8.28% 12.31% 76.74%
Comparing the results between the second and third rows in Table 9, we find that PCA facilitates FC in learning more useful information from raw level circumferences, leading to a 4.28g decrease in MAE and 18.60% increase in accuracy. Further comparison between the third, fourth and last rows indicates that using either PCA or RNN alone does not yield optimal results. The incorporation of both, as implemented in our final design, presents the most significant improvement in performance. The fifth row shows the performance results when the RNN is replaced with GRU in our network. Despite the GRU’s advanced capabilities in capturing long-term relationships from sequences, this modification does not bring more improvement in this task. This failure could potentially be due to the fact that GRU contains more trainable parameters than the conventional RNN, thus resulting in reduced learning efficiency, particularly when dealing with small data size. These findings affirm the effectiveness of our proposed hybrid structure for 3D body feature extraction.
6 Conclusion
In this study, we examined the potential of using 3D body scans with few basic demographic information for prenatal care for future telehealth application and explored how to use this modality efficiently. We proposed a novel neural network which incorporates a supervised learning stream and an unsupervised learning stream for extracting features from sampled abdominal level circumferences. The supervised stream leverages RNN units to extract sequential information from level circumferences, while the unsupervised stream employs the PCA to captures global descriptors of abdominal body shape. We applied this algorithm to predict the risk of preterm labor, GDM, GH, the likelihood of cesarean section, and to estimate current fetal weight and MVP. The results indicate that maternal 3D body shape, captured during 18-24 gestational weeks, is effective in predicting GDM, GH, and preterm labor, as well as estimating fetal weight. In comparison to other well-performing machine learning algorithms, our proposed algorithm demonstrates superior performances in these tasks.
Nevertheless, there is room for future improvement. Firstly, our current dataset is small and only contains scans obtained in patients’ second trimester of pregnancy, potentially limiting the model’s generalizability to estimate fetal weight in other trimesters, since fetal growth trends may vary across different stages. Ideally, with the inclusion of actual birth weights and 3D body scans closer to delivery, a more accurate model may be developed guided by neonatal weights. We may also explore the potential of using longitudinal data for better estimation. Secondly, our current analysis focused solely on the abdominal region, overlooking other body parts. Given the research [17, 47, 48, 49] suggesting correlations between appendicular body shape and pregnancy outcomes, future studies could investigate methods using the entire body shape to get more accurate estimations and predictions. To enable the practical application of 3D body scans in telehealth, we will also conduct further experiments using smartphone-scanned 3D models.
[…]
Conflict of interest
The authors have no conflicts of interest to disclose.
[…]
Data availability statement
Data may be made available based on the nature of the request.
[…]
Funding statement
Research reported in this publication was supported by the National Institute of Diabetes and Digestive and Kidney Diseases of the National Institutes of Health under Award Number R01DK129809. The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health.
Author Biographies
Ruting Cheng is a Ph.D. student at George Washington University. She received her M.S. degree from the University of Science and Technology Beijing in 2022. Her research interests include computer vision and machine learning in medicine.
Yijiang Zheng is a Ph.D. student at George Washington University, where he earned his master’s degree in computer science in 2020. His research focuses on 3D computer vision applications within the medical field.
Boyuan Feng is a Ph.D. student in Computer Science at George Washington University and holds a B.S. from Huazhong University of Science and Technology. Her research focuses on machine learning and computer vision.
Chuhui Qiu is a graduate student at the George Washington University, where he received his M.S. degree. His field of study is machine learning, computer vision, and algorithm design.
Zhuoxin Long is currently a Ph.D. student at the George Washington University, where she earned her master’s degree in Biostatistics. Her research interests include causal inference and individualized treatment regimes.
Joaquin A. Calderon earned his MD from Universidad del Norte in Barranquilla, Colombia in 2018. He joined the GW Medical Faculty Associates in 2022 as a Research Coordinator in Obstetrics and Gynecology. He is interested in maternal-fetal medicine.
Xiaoke Zhang is currently an Associate Professor in the Department of Statistics at George Washington University. His research interests include longitudinal data analysis, statistical learning, causal inference and applied statistics.
Jaclyn M. Phillips, MD is a Maternal Fetal Medicine specialist and Assistant Professor of Obstetrics & Gynecology. Her research interests include postpartum hemorrhage, severe maternal morbidity, peripartum blood management, and risk prediction.
James K. Hahn is a Professor in both the Department of Computer Science, and School of Medicine and Health Sciences at the George Washington University. His areas of interests are: medical simulation, machine learning and medical visualization.